in the standard (x,y) coordinate plane, what is the slope of the line 11x +6y =3
Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

11x + 6y = 3
6y = -11x + 3 {subtracted 11x from each side}
y = (-11/6)x + (1/2) {divided each side by 6}

slope = -11/6

For more help from me, visit: www.algebrahouse.com